OpenStudy (anonymous):

What is √(14q) x 2 √(4q) ? Can you show me how you got it

OpenStudy (anonymous):

@jim_thompson5910 help?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{14q}\times 2\sqrt{4q}\] \[\large 2\sqrt{14q}\times\sqrt{4q}\] \[\large 2\sqrt{14q*4q}\] \[\large 2\sqrt{56q^2}\] \[\large 2\sqrt{14*4q^2}\] \[\large 2\sqrt{4q^2*14}\] \[\large 2\sqrt{4q^2}*\sqrt{14}\] I'll let you finish up

OpenStudy (anonymous):

\[2\sqrt{56q ^{2}}\] ?

jimthompson5910 (jim_thompson5910):

i already have that step (see line 4)

jimthompson5910 (jim_thompson5910):

but notice how I moved on from it (so that means that it's not completely simplified there)

OpenStudy (anonymous):

I dont understand how to do more then that. sorry

jimthompson5910 (jim_thompson5910):

how would you simplify \[\large \sqrt{4q^2}\]

OpenStudy (anonymous):

2q?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{14q}\times 2\sqrt{4q}\] \[\large 2\sqrt{14q}\times\sqrt{4q}\] \[\large 2\sqrt{14q*4q}\] \[\large 2\sqrt{56q^2}\] \[\large 2\sqrt{14*4q^2}\] \[\large 2\sqrt{4q^2*14}\] \[\large 2\sqrt{4q^2}*\sqrt{14}\] \[\large 2*2q*\sqrt{14}\] \[\large 4q*\sqrt{14}\] and that's all we can do

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np